ABC Company has variable operating cost of $15 per unit and sell price of 22 per unit if company’s operating break-even point is 3000 unit:

  1. What is the fixed operating cost?

  2. If company produce 500 units what is company’s EBIT?

  3. Calculate the degree of operating leverage when company make 4000 units.
Answer #1

1)

Fixed Operating cost = contribution * Break even point

Contribution = selling price - Variable cost

= 22 - 15

= 7

Fixed Operating cost = 7*3000 = $21,000

2)

EBIT = (sales * contribution) - fixed costs

= (500 * 7) - 21,000

= -$17,500

3)

Degree of operating leverage = contribution / EBIT

contribution = 4000*7 = 28000

EBIT = contribution - fixed costs

= 28,000 - 21,000 = 7,000

DOL = 28,000 / 7,000

= 4